Classify each system as consistent or inconsistent:
{2x+3(y-2)=0
{-3y=2(x-4)

Answer by stanbon(75887)

Classify each system as consistent or inconsistent:
{2x+3(y-2)=0
{-3y=2(x-4)
===================
Simplify:
2x +3y = 6
2x +3y = 8
=================
The x value and the y value in
the two equations cannot be the
same.
The system is inconsistent.
=================================
Cheers,
Stan H.